OpenStudy (anonymous):

(4x – 3) (2x^2 – 7x + 1)

OpenStudy (anonymous):

Will it be 8x^3 - 28x + 6 -6x^2 + 21 - 3

OpenStudy (anonymous):

@mathslover

thomaster (thomaster):

4x*2x^2= 8x^3 4x*-7x -28x^2 4x*1 = 4x -3*2x^2 = -6x^2 -3*-7x= 21x -3*1 =-3 8x^3-28x^2+4x-6x^2+21x-3 8x^3 - 34x^2 - 17x - 3

thomaster (thomaster):

wait i made a mistake its not -17x its +25x \(8x^3−34x^2+25x−3\)